Update at 20:54:47( BST ) : The issue has now been resolved with this server.
22nd April 2010, at 6:59PM
22nd April 2010, at 7:54PM
We applied a potential fix to the FTP lock problem late yesterday evening. If your FTP lock is taking longer than 10-15 minutes to release, please contact support to notify them of the domain name in question.
21st April 2010, at 7:18PM
22nd April 2010, at 5:13PM
The load averages have dropped significantly since yesterday evening and the server appears to be running well.
We will keep a close eye on this server throughout the day.
21st April 2010, at 6:22PM
22nd April 2010, at 5:13PM
Update at 18:43:03( BST ) : The issue has now been resolved with this server.
18th April 2010, at 6:42PM
18th April 2010, at 5:43PM
19/04/10 @ 22:10 We have managed to control the level of traffic affecting the network, services should be back up to full speed.
19/04/10 @ 22:00 We are looking at a repeat of this issue at this moment. Some services will be temporarily slowed.
--
We've seen quite a bit of network congestion this afternoon and believe to have found the malicious content that was causing said issues.
We'll continue to monitor this and shall contact the user in question.
18th April 2010, at 6:39PM
20th April 2010, at 8:37PM
Mail queues are now sufficiently low on the majority of servers to warrant lowering this issue to the 'monitoring' status. We will of course continue to closely monitor our SMTP servers for now.
16th April 2010, at 10:58AM
20th April 2010, at 8:37PM
All checks completed and we're now happy to resolve this status.
15th April 2010, at 4:45PM
15th April 2010, at 6:04PM
Update at 13:54:18( BST ) : The issue has now been resolved with this server.
14th April 2010, at 1:37PM
14th April 2010, at 12:54PM
This status is being resolved as VPS2 has performed satisfactorily over the last week. Nevertheless, we have taken the decision to begin migrating VPS' away from some of our older hosts and will contact any affected customers in due course. Any customers that wish to be migrated early are welcome to contact us ahead of time.
14th April 2010, at 11:35AM
18th April 2010, at 5:41PM
Update at 09:15:40( BST ) : The issue has now been resolved with this server.
14th April 2010, at 9:10AM
14th April 2010, at 8:15AM